# CurvFed: Curvature-Aligned Federated Learning for Fairness without Demographics

### v6

@@ 1. Introduction
The integration of advanced machine learning (ML) with ubiquitous sensing (li2021deep, ; puccinelli2005wireless, ) has enabled everyday devices like smartphones and wearables to collect rich data for human-centric applications, from gesture recognition (zhang2021WIDAR3, ) to stress detection (vos2023generalizable, ). Concurrently, edge and distributed computing—especially federated learning (FL)—have emerged as key frameworks for decentralized model training (zhou2019edge, ; mcmahan2017communication, ). In FL, each client (e.g., device or organization) retains local data and trains models privately (li2024comprehensive, ), reducing centralization risks (arcuri2010centralization, ; brynjolfsson2023big, ) and aligning with privacy regulations like GDPR, CCPA, and HIPAA (zhang2020fairfl, ; abay2020mitigating, ). This makes FL ideal for human-sensing applications like health monitoring and activity recognition (epstein2020mapping, ; kim2016timeaware, ; shaik2022fedstack, ).
Notably, recent studies have begun exploring fairness in human-sensing federated learning (FL) (yfantidou2023beyond, ; xu2023globem, ; zhang2023framework, ). However, state-of-the-art FL fairness approaches face a key challenge due to the trade-off between fairness and privacy (chen2023privacy, ), often requiring access to sensitive information at the client or server level, which contradicts FL’s privacy-preserving nature (zeng2021improving, ; papadaki2022minimax, ; djebrouni2024bias, ; zhang2020fairfl, ; ezzeldin2023fairfed, ). This underscores the need for fairness-enhancing approaches that do not rely on sensitive or bias-inducing attributes. In human-sensing FL, this is especially important, as unlabeled factors can bias models —e.g., building materials can skew Wi-Fi CSI-based activity detection (lee2019effects, ).
These challenges motivate the pursuit of Fairness without Demographics (FWD) (lahoti2020fairness, ) -refers to achieving equitable model performance across users or groups without relying on sensitive or demographic attributes such as sex, age, materials, among others, in FL, particularly since most human-sensing datasets lack labeled sensitive attributes (yfantidou2023beyond, ; mehrabi2021survey, ; zhang2023framework, ). This also aligns with the 2025 Executive Order’s (house2025removing, ) vision for bias-resistant, innovation-driven AI by promoting fairness through optimization principles—without relying on demographic or group attributes—supporting equitable systems free from engineered social agendas. While FWD has been studied in centralized AI training (lahoti2020fairness, ; hashimoto2018fairness, ; chai2022fairness, ), and only one FWD in FL work named FedSRCVaR (papadaki2024federated, ) exists, FWD’s integration into privacy-preserving decentralized FL remains challenging in human-centric FL, as outlined below:
@@
    Centralized FWD methods (Section 2.2) depend on access to global data distributions to identify biases and adjust decision boundaries, which is unavailable in privacy-preserving FL where clients only access local data (abbas2023context, ), making fairness enforcement more difficult.
@@
    Although FedSRCVaR (papadaki2024federated, ) does not require sensitive attribute labels in data, it relies on the knowledge of the bounds for sensitive or worst-case group sizes, making it not entirely agnostic to sensitive attribute knowledge (see Section 2.2).
@@
    Traditional group fairness methods—including FWD approaches (Section 2.2)—generally assume that all sensitive groups are represented during optimization (rieke2020future, ). However, this assumption is impractical in human-sensing FL, where each client often corresponds to a single user or a narrow demographic segment, resulting in an incomplete representation of population-level groups.
@@
    In human-centric sensing, multiple latent biases (e.g., gender, body type, sensor placement) coexist (jui2024fairness, ), and mitigating one bias may inadvertently worsen another (djebrouni2024bias, ).
@@
CurvFed utilizes the Fisher Information Matrix (FIM) as a fast-computable proxy for loss-landscape sharpness (lee2022masking, ; thomas_interplay_2020, ), making it well-suited for resource-constrained edge devices (as FL clients), where it regularizes locally trained models to enhance fairness. Importantly, it introduces a sharpness-aware aggregation scheme, which aligns loss-landscape curvature across all clients by leveraging client-wise FIM data. This addresses the above-discussed third challenge, since FIM can be computed for both single-user and multi-user clients, allowing CurvFed to operate without requiring full representation of sensitive attributes within each client.
@@
  • Superior fairness–accuracy tradeoff: Existing FL methods often struggle with the inherent tension between fairness and accuracy (ezzeldin2023fairfed, ; gu2022privacy, ). CurvFed mitigates this by promoting flatter loss landscapes during client training, which inherently enhances generalizability (dauphin2024neglected, ; jastrzkebski2018relation, ) and model inference confidence by pushing samples away from decision boundaries (tran2022pruning, ). It further optimizes both curvature smoothness and accuracy during aggregation, yielding a better balance of fairness and performance (see Sections 4.1 and 5.2, RQ-1).
@@
  • Improved multi-attribute fairness simultaneously: Human-sensing data involves multiple sensitive attributes (zhang2020fairfl, ; ezzeldin2023fairfed, ), but fair FL methods struggle to address them simultaneously (zhang2020fairfl, ; cui2021addressing, ). CurvFed promotes flatter client loss landscapes, improving fairness across all attributes without relying on attribute labels (see Section 5.2, RQ-4).
  • Real-world feasibility and sensitivity analysis: We validate CurvFed’s practicality through extensive evaluations across data volume, client sampling, communication cost, resource usage, and runtime on diverse platforms. Additionally, a real-world testbed evaluation on a heterogeneous testbed of six edge devices confirms its effectiveness under non-simulated conditions (see Sections 5.3 and 5.4).
We demonstrate the effectiveness of CurvFed across three real-world human-sensing datasets, each representing different sensing modalities, applications, and sensitive attributes, showing CurvFed’s effectiveness in FWD. We adopt off-the-shelf models previously published for these datasets (zhang2021WIDAR3, ; xiao2024reading, ; chereshnev2018hugadb, ) for FL setting. However, CurvFed is architecture-agnostic. The framework is not tied to specific model choices and can be readily extended to diverse use cases and architectures. The rest of the paper is organized as follows: Section 2 reviews prior work; Section 3 presents the theoretical foundation and assumptions of CurvFed; Section 3.1 defines the problem, gaps, and challenges. Section 4 details the proposed method, Section 5 presents in-depth results, Section 6 presents empirical verifications of our propositions, and Section 7 discusses broader impacts and limitations.
@@ 2. Background and Related works
Fairness has emerged as a critical concern in machine learning (ML) research, with early efforts focusing on pre-processing strategies, such as removing sensitive attributes from training data (dwork2012fairness, ). However, such approaches often compromise model performance (zhang2020fairfl, ), prompting a shift toward in-processing methods that embed fairness constraints into model optimization (zhang2019faht, ). While effective in centralized ML, these methods face new challenges in federated learning (FL), where privacy constraints limit access to sensitive attributes (pessach2022review, ; djebrouni2024bias, ). This section discusses the recent fairness approaches in FL, highlights common evaluation metrics, and motivates the need for Fairness Without Demographics (FWD).
@@ 2.1. Fairness in Federated Learning
Bias in FL systems can arise in multiple ways (li2023hierarchical, ; zeng2021improving, ; djebrouni2024bias, ; cui2021addressing, ; zhao2018federated, ), but a primary concern is performance-based group fairness, which aims to reduce disparities in model performance across demographic groups (castelnovo2022clarification, ). This is distinct from contribution fairness, which addresses disparities from client participation or data imbalance (e.g., non-IID data) (li2023hierarchical, ; zhao2018federated, ), and is beyond the scope of this work. In this study, we align with performance-based group fairness approaches and focus on ensuring that the global model does not disproportionately underperform for any groups, a challenge exacerbated by the decentralized nature of data in FL.
@@
Client-Side and Hybrid Approaches: Several methods enforce fairness during local training. For example, FedFB applies fair batch sampling on the client side (zeng2021improving, ), while others use min-max optimization to minimize worst-case group loss (mohri2019agnostic, ; du2021fairness, ; papadaki2022minimax, ). However, these approaches typically require access to sensitive attribute labels, making them unsuitable for Fairness Without Demographics (FWD). Notably, AFL (mohri2019agnostic, ) targets fairness across clients, rather than demographic groups, whereas others assume the availability of knowledge of sensitive group labels and distributions.
Recent works that implicitly flatten the loss landscape offer a closer fit to our setting, as they can promote fairness without accessing sensitive data. For example, (caldarola2022improving, ) applies Sharpness-Aware Minimization (SAM) at the client and Stochastic Weight Averaging (SWA) at the server. These methods, which do not rely on demographic attributes, serve as relevant baselines for our evaluations. Our approach, CurvFed, explicitly leverages sharpness cues and follows a hybrid framework leveraging sharpness information from the loss surface (a manner distinct from SAM) and adapting SWA in one of the modules. Our ablation study in Section 5.5 also demonstrates the impact of such adaptation in CurvFed.
Server-Side Approaches: Some strategies address fairness during the server’s aggregation phase. FairFed (ezzeldin2023fairfed, ) adjusts model aggregation weights based on local fairness evaluations performed by clients. Similarly, Astral (djebrouni2024bias, ) uses an evolutionary algorithm to select clients for aggregation to promote fairness. While these methods maintain data decentralization, they still require clients to compute fairness metrics using sensitive group information, creating potential privacy risks and conflicting with the goals of FWD.
@@ 2.2. Fairness Without Demographics (FWD)
The challenge of achieving fairness without access to sensitive attributes has motivated the field of Fairness Without Demographics (FWD). In centralized ML, this has been explored through techniques like adversarial re-weighting to boost the importance of underperforming instances (lahoti2020fairness, ). However, in FL, each client operates with its own training instances, making it challenging to identify such instances and apply this approach directly. Another approach, knowledge distillation (KD), was used by (chai2022fairness, ) to transfer knowledge from a complex teacher to a student model to improve fairness without accessing sensitive information. We adapt this KD concept to FL as a baseline. After initial few rounds of initial training via the FedAvg (mcmahan2017communication, ) protocol, each client subsequently performs KD using the global model as the teacher, training a local student model on its own data with soft labels from the teacher before sharing their local model for aggregation.
To our knowledge, the only existing work applying FWD directly to FL is FedSRCVaR (papadaki2024federated, ). While both our work and FedSRCVaR target fairness without sensitive attributes, their problem formulations differ fundamentally. FedSRCVaR adopts a Rawlsian approach (lahoti2020fairness, ), optimizing for a worst-performing subset of individuals defined by a tunable threshold, $\rho$. Selecting an appropriate $\rho$ is non-trivial in heterogeneous FL settings and requires prior knowledge or extensive tuning (guo2021towards, ). Furthermore, its evaluation assumes each client represents a single sensitive group. In contrast, our setting assumes clients may have mixed or sparse group representations. Despite a lack of public implementation details, we adapted FedSRCVaR to our setting and report it as a key baseline for a comprehensive evaluation. Our method, CurvFed, differs by not imposing constraints on group size, offering a more flexible FWD approach.
@@ 2.3. Evaluation Metrics
We adopt established metrics to assess fairness and its trade-off with accuracy in FL (ezzeldin2023fairfed, ; djebrouni2022towards, ; xu2023fairadabn, ; chiu2024achieve, ; li2024bffn, ).
Fairness: We use the Equal Opportunity (EO) gap (ezzeldin2023fairfed, ; djebrouni2022towards, ), defined as:
(1) $\Delta_{\text{EO}}=P(\hat{Y}=1|S=1,Y=1)-P(\hat{Y}=1|S=0,Y=1)$
A lower $\Delta_{\text{EO}}$ indicates greater fairness. Sensitive attributes are used only during post-training evaluation.
Fairness–Accuracy Trade-off: To evaluate balance, we use the FATE score (xu2023fairadabn, ; chiu2024achieve, ; li2024bffn, ), which compares normalized changes in accuracy and fairness relative to a benign baseline:
@@
Here, $m$ and $b$ denote the fair model and the baseline (FedAvg (mcmahan2017communication, )). A higher FATE score reflects a better fairness–accuracy trade-off.

### v7

@@ 1. Introduction
The integration of advanced machine learning (ML) with ubiquitous sensing (Li et al., 2021a; Puccinelli and Haenggi, 2005) has enabled everyday devices like smartphones and wearables to collect rich data for human-centric applications, from gesture recognition (Zhang et al., 2021) to stress detection (Vos et al., 2023). Concurrently, edge and distributed computing—especially federated learning (FL)—have emerged as key frameworks for decentralized model training (Zhou et al., 2019; McMahan et al., 2017). In FL, each client (e.g., device or organization) retains local data and trains models privately (Li et al., 2024a), reducing centralization risks (Arcuri and Dari-Mattiacci, 2010) and aligning with privacy regulations like GDPR, CCPA, and HIPAA (Zhang et al., 2020; Abay et al., 2020). This makes FL ideal for human-sensing applications like health monitoring and activity recognition (Epstein et al., 2020; Kim et al., 2016; Shaik et al., 2022).
Notably, recent studies have begun exploring fairness in human-sensing federated learning (FL) (Yfantidou et al., 2023a; Xu et al., 2023a; Zhang et al., 2023b). However, state-of-the-art FL fairness approaches face a key challenge due to the trade-off between fairness and privacy (Chen et al., 2023), often requiring access to sensitive information at the client or server level, which contradicts FL’s privacy-preserving nature (Zeng et al., 2021; Papadaki et al., 2022; Djebrouni et al., 2024; Zhang et al., 2020; Ezzeldin et al., 2023). This underscores the need for fairness-enhancing approaches that do not rely on sensitive or bias-inducing attributes. In human-sensing FL, this is especially important, as unlabeled factors can bias models —e.g., building materials can skew Wi-Fi CSI-based activity detection (Lee et al., 2019).
These challenges motivate the pursuit of Fairness without Demographics (FWD) (Lahoti et al., 2020) -refers to achieving equitable model performance across users or groups without relying on sensitive or demographic attributes such as sex, age, materials, among others, in FL, particularly since most human-sensing datasets lack labeled sensitive attributes (Yfantidou et al., 2023a; Mehrabi et al., 2021; Zhang et al., 2023b). This also aligns with the 2025 Executive Order’s (House, 2025) vision for bias-resistant, innovation-driven AI by promoting fairness through optimization principles—without relying on demographic or group attributes—supporting equitable systems free from engineered social agendas. While FWD has been studied in centralized AI training (Lahoti et al., 2020; Hashimoto et al., 2018; Chai et al., 2022), and only one FWD in FL work named FedSRCVaR (Papadaki et al., 2024) exists, FWD’s integration into privacy-preserving decentralized FL remains challenging in human-centric FL, as outlined below:
@@
    Centralized FWD methods (Section 2.2) depend on access to global data distributions to identify biases and adjust decision boundaries, which is unavailable in privacy-preserving FL where clients only access local data (Abbas et al., 2023), making fairness enforcement more difficult.
@@
    Although FedSRCVaR (Papadaki et al., 2024) does not require sensitive attribute labels in data, it relies on the knowledge of the bounds for sensitive or worst-case group sizes, making it not entirely agnostic to sensitive attribute knowledge (see Section 2.2).
@@
    Traditional group fairness methods—including FWD approaches (Section 2.2)—generally assume that all sensitive groups are represented during optimization (Rieke et al., 2020). However, this assumption is impractical in human-sensing FL, where each client often corresponds to a single user or a narrow demographic segment, resulting in an incomplete representation of population-level groups.
@@
    In human-centric sensing, multiple latent biases (e.g., gender, body type, sensor placement) coexist (Jui and Rivas, 2024), and mitigating one bias may inadvertently worsen another (Djebrouni et al., 2024).
@@
CurvFed utilizes the Fisher Information Matrix (FIM) as a fast-computable proxy for loss-landscape sharpness (Lee et al., 2022; Thomas et al., 2020), making it well-suited for resource-constrained edge devices (as FL clients), where it regularizes locally trained models to enhance fairness. Importantly, it introduces a sharpness-aware aggregation scheme, which aligns loss-landscape curvature across all clients by leveraging client-wise FIM data. This addresses the above-discussed third challenge, since FIM can be computed for both single-user and multi-user clients, allowing CurvFed to operate without requiring full representation of sensitive attributes within each client.
@@
  • Superior fairness–accuracy tradeoff: Existing FL methods often struggle with the inherent tension between fairness and accuracy (Ezzeldin et al., 2023; Gu et al., 2022). CurvFed mitigates this by promoting flatter loss landscapes during client training, which inherently enhances generalizability (Dauphin et al., 2024; Jastrzębski et al., 2018) and model inference confidence by pushing samples away from decision boundaries (Tran et al., 2022). It further optimizes both curvature smoothness and accuracy during aggregation, yielding a better balance of fairness and performance (see Sections 4.1 and 5.3, RQ-1).
@@
  • Improved multi-attribute fairness simultaneously: Human-sensing data involves multiple sensitive attributes (Zhang et al., 2020; Ezzeldin et al., 2023), but fair FL methods struggle to address them simultaneously (Zhang et al., 2020; Cui et al., 2021). CurvFed promotes flatter client loss landscapes, improving fairness across all attributes without relying on attribute labels (see Section 5.3, RQ-4).
  • Real-world feasibility and sensitivity analysis: We validate CurvFed’s practicality through extensive evaluations across data volume, client sampling, communication cost, resource usage, and runtime on diverse platforms. Additionally, a real-world testbed evaluation on a heterogeneous testbed of six edge devices confirms its effectiveness under non-simulated conditions (see Sections 6 and 6.2).
We demonstrate the effectiveness of CurvFed across three real-world human-sensing datasets, each representing different sensing modalities, applications, and sensitive attributes, showing CurvFed’s effectiveness in FWD. We adopt off-the-shelf models previously published for these datasets (Zhang et al., 2021; Xiao et al., 2024; Chereshnev and Kertész-Farkas, 2018) for FL setting. However, CurvFed is architecture-agnostic. The framework is not tied to specific model choices and can be readily extended to diverse use cases and architectures. The rest of the paper is organized as follows: Section 2 reviews prior work; Section 3 presents the theoretical foundation and assumptions of CurvFed; Section 3.1 defines the problem, gaps, and challenges. Section 4 details the proposed method, Section 5 presents in-depth results, Section 7 presents empirical verifications of our propositions, and Section 8 discusses broader impacts and limitations.
@@ 2. Background and Related works
Fairness has emerged as a critical concern in machine learning (ML) research, with early efforts focusing on pre-processing strategies, such as removing sensitive attributes from training data (Dwork et al., 2012). However, such approaches often compromise model performance (Zhang et al., 2020), prompting a shift toward in-processing methods that embed fairness constraints into model optimization (Zhang and Ntoutsi, 2019). While effective in centralized ML, these methods face new challenges in federated learning (FL), where privacy constraints limit access to sensitive attributes (Pessach and Shmueli, 2022; Djebrouni et al., 2024). This section discusses the recent fairness approaches in FL, highlights common evaluation metrics, and motivates the need for Fairness Without Demographics (FWD).
@@ 2.1. Fairness in Federated Learning
Bias in FL systems can arise in multiple ways (Li et al., 2023; Zeng et al., 2021; Djebrouni et al., 2024; Cui et al., 2021; Zhao et al., 2018), but a primary concern is performance-based group fairness, which aims to reduce disparities in model performance across demographic groups (Castelnovo et al., 2022). This is distinct from contribution fairness, which addresses disparities from client participation or data imbalance (e.g., non-IID data) (Li et al., 2023; Zhao et al., 2018), and is beyond the scope of this work. In this study, we align with performance-based group fairness approaches and focus on ensuring that the global model does not disproportionately underperform for any groups, a challenge exacerbated by the decentralized nature of data in FL.
@@
Client-Side and Hybrid Approaches: Several methods enforce fairness during local training. For example, FedFB applies fair batch sampling on the client side (Zeng et al., 2021), while others use min-max optimization to minimize worst-case group loss (Mohri et al., 2019; Du et al., 2021; Papadaki et al., 2022). However, these approaches typically require access to sensitive attribute labels, making them unsuitable for Fairness Without Demographics (FWD). Notably, AFL (Mohri et al., 2019) targets fairness across clients, rather than demographic groups, whereas others assume the availability of knowledge of sensitive group labels and distributions.
Recent works that implicitly flatten the loss landscape offer a closer fit to our setting, as they can promote fairness without accessing sensitive data. For example, (Caldarola et al., 2022) applies Sharpness-Aware Minimization (SAM) at the client and Stochastic Weight Averaging (SWA) at the server. These methods, which do not rely on demographic attributes, serve as relevant baselines for our evaluations. Our approach, CurvFed, explicitly leverages sharpness cues and follows a hybrid framework leveraging sharpness information from the loss surface (a manner distinct from SAM) and adapting SWA in one of the modules. Our ablation study in Section 6.3 also demonstrates the impact of such adaptation in CurvFed.
Server-Side Approaches: Some strategies address fairness during the server’s aggregation phase. FairFed (Ezzeldin et al., 2023) adjusts model aggregation weights based on local fairness evaluations performed by clients. Similarly, Astral (Djebrouni et al., 2024) uses an evolutionary algorithm to select clients for aggregation to promote fairness. While these methods maintain data decentralization, they still require clients to compute fairness metrics using sensitive group information, creating potential privacy risks and conflicting with the goals of FWD.
@@ 2.2. Fairness Without Demographics (FWD)
The challenge of achieving fairness without access to sensitive attributes has motivated the field of Fairness Without Demographics (FWD). In centralized ML, this has been explored through techniques like adversarial re-weighting to boost the importance of underperforming instances (Lahoti et al., 2020). However, in FL, each client operates with its own training instances, making it challenging to identify such instances and apply this approach directly. Another approach, knowledge distillation (KD), was used by (Chai et al., 2022) to transfer knowledge from a complex teacher to a student model to improve fairness without accessing sensitive information. We adapt this KD concept to FL as a baseline. After initial few rounds of initial training via the FedAvg (McMahan et al., 2017) protocol, each client subsequently performs KD using the global model as the teacher, training a local student model on its own data with soft labels from the teacher before sharing their local model for aggregation.
To our knowledge, the only existing work applying FWD directly to FL is FedSRCVaR (Papadaki et al., 2024). While both our work and FedSRCVaR target fairness without sensitive attributes, their problem formulations differ fundamentally. FedSRCVaR adopts a Rawlsian approach (Lahoti et al., 2020), optimizing for a worst-performing subset of individuals defined by a tunable threshold, $\rho$. Selecting an appropriate $\rho$ is non-trivial in heterogeneous FL settings and requires prior knowledge or extensive tuning (Guo et al., 2021). Furthermore, its evaluation assumes each client represents a single sensitive group. In contrast, our setting assumes clients may have mixed or sparse group representations. Despite a lack of public implementation details, we adapted FedSRCVaR to our setting and report it as a key baseline for a comprehensive evaluation. Our method, CurvFed, differs by not imposing constraints on group size, offering a more flexible FWD approach.
2.3. A Loss landscape Primer
The loss landscape refers to the geometry of the objective function (or loss) mapped over the high-dimensional space of model parameters. The geometric properties of this landscape, particularly around the local minima found during training, are critical indicators of model performance (Li et al., 2018).
Curvature and Sharpness: Curvature describes the rate at which the loss function changes as model parameters are perturbed(Li et al., 2018). Mathematically, this is characterized by the Hessian matrix (the matrix of second-order derivatives) of the loss function(Tran et al., 2022; Sankar et al., 2021). Regions with high curvature are often referred to as sharp minima, where small changes in parameters lead to significant increases in loss. Conversely, flat minima are regions with low curvature, where the loss remains relatively stable despite parameter perturbations.
Connection to Model overfitting: Prior research has established a strong link between the geometry of the loss landscape and model generalization. "Flat" minima are generally associated with better generalization to unseen data, as they indicate a solution that is robust to shifts in the data distribution or noisy parameter updates (Hochreiter and Schmidhuber, 1997; Keskar et al., 2017). In contrast, sharp minima often lead to model overfitting (Wen et al., 2018; Keskar et al., 2017).
Measuring Curvature: While the Hessian provides a precise measure of curvature, computing it is computationally prohibitive for deep neural networks (Singh and Alistarh, 2020). In this work, we utilize the top eigenvalue of the Fisher Information Matrix (FIM) as a computationally efficient proxy for the Hessian’s top eigenvalue (and thus, the sharpness) (Karakida et al., 2019; Liang et al., 2019). This metric allows us to quantify and align the "flatness" of models across federated clients without incurring the heavy cost of full Hessian computation.
@@ 2.4. Evaluation Metrics
We adopt established metrics to assess fairness and its trade-off with accuracy in FL (Ezzeldin et al., 2023; Djebrouni, 2022; Xu et al., 2023b; Chiu et al., 2024; Li et al., 2024b).
Fairness: We use the Equal Opportunity (EO) gap (Ezzeldin et al., 2023; Djebrouni, 2022), defined as:
(1) $\Delta_{\text{EO}}=P(\hat{Y}=1\mid S=1,Y=1)-P(\hat{Y}=1\mid S=0,Y=1)$
where $\hat{Y}$ denotes the model’s predicted label, $Y$ the ground-truth label, and $S\in\{0,1\}$ the binary sensitive attribute (e.g., sex or hand). The EO gap measures disparities in true positive rates across sensitive groups; a lower $\Delta_{\text{EO}}$ indicates greater fairness. Sensitive attributes are used only during post-training evaluation.
Fairness–Accuracy Trade-off: To evaluate balance, we use the FATE score (Xu et al., 2023b; Chiu et al., 2024; Li et al., 2024b), defined as:
@@
where ACC denotes classification accuracy, EO the Equal Opportunity gap, subscript $m$ refers to the fair model, and subscript $b$ denotes the baseline method (FedAvg (McMahan et al., 2017)). The FATE score jointly captures relative accuracy gains and fairness improvements; higher values indicate a more favorable fairness–accuracy trade-off.